6x – 0.37 = 5(x – 0.28)

Answer by pwac(253)

6x-0.37=5(x-0.28) remove bracket by multiplying by 5
6x-0.37=5x-1.4 take x's to left,number to right changing signs
6x-5x=-1.4+0.37
x=-1.03
Pete